$SPAD/src/input manuel.input

Manuel Bronstein

December 3, 2016

Abstract

This integral from Manuel Bronstein was posted on comp.soft-sys.math.maple as one that Axiom handled with a simple result but generated long output on Mathematica and Maple.

Contents

2

— * —

)set break resume
)spool manuel.output
)set message test on
)set message auto off
)clear all
--S 1 of 3
g:=x/sqrt(x^4+10*x^2-96*x-71)
--R
--R                   x
--R   (1)  -----------------------
--R         +--------------------+
--R         | 4      2
--R        \|x  + 10x  - 96x - 71
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 3
f:=integrate(g,x)
--R
--R   (2)
--R   -
--R        log
--R                                                    +--------------------+
--R               6      4      3      2               | 4      2                8
--R             (x  + 15x  - 80x  + 27x  - 528x + 781)\|x  + 10x  - 96x - 71  - x
--R           +
--R                  6       5      4        3        2
--R             - 20x  + 128x  - 54x  + 1408x  - 3124x  - 10001
--R     /
--R        8
--R                                         Type: Union(Expression(Integer),...)
--E 2
--S 3 of 3
differentiate(f,x)-g
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 3
)spool
)lisp (bye)

———-

References^{[1] nothing}

4